A credit card bill for $559 was due on September 14. Purchases of $289 were made on September 19, and $18 was charged on September 28. A payment of $250 was made on September 25. The annual interest on the average daily balance is 21.5%. Find the finance charge due (in dollars) on the October 14 bill. (Use 365 for the number of days in a year. Enter a number. Round your answer to the nearest cent.)
Answer #1

it is assumed that bill due on 14 September is not paid

calculation of Finance charge due on October 14

(+) $ 559 due on sep 14, $ 559 * (21.5*31) / (100 * 365) = $ 10.20

(+) purchase $ 289 on sep 19, $ 289 * (21.5*26 ) / (100 * 365) = $ 4.42

(+) finance charge on sep 28, $ 18 *(21.5*17 ) / (100 * 365) = $ 0.1802

(-) Repayment on 25 sep , $ 250 * (21.5*20 ) / (100 * 365) = (2.945)

Total finance charges from 14 sep to 14 oct = $ 11.855

Amount due on 14 October = $559+$289+$18+$11.855- $250

= $ 627.855